OpenStudy (anonymous):

-4(t-3)-(t+6)=9 i got t=-3 but when i check it dont check out

OpenStudy (lgbasallote):

let's try \[\implies -4(-3 - 3) - (-3 + 6) = 9\] \[\implies -4(-6) - (3) = 9\] \[\implies 24 - 3 = 9\] \[\implies 21 \ne 9\]

OpenStudy (lgbasallote):

let's try resolving \[-4(t - 3) - (t + 6) = 9\] distribute.. \[\implies -4t + 12 - t - 6 = 9\] combine like terms... \[\implies -5t + 6 = 9\] subtract 6 from both sides... \[\implies -5t = 3\] do you see it now?

OpenStudy (anonymous):

oh i had subtracted 6 from 6 and added 6 to the 9 the first time

OpenStudy (lgbasallote):

oh lol. well now you know

OpenStudy (anonymous):

so t=-0.6

OpenStudy (lgbasallote):

right. but you might want to put that in fraction form...just in case

OpenStudy (anonymous):

thanks

OpenStudy (lgbasallote):

welcome